# Graph-based Square-Root Estimation for Sparse Linear Regression

### v1

@@ 4.1 Simulation data
Table 2: Performance comparison of estimation, prediction and model selection for Example 1&2 with t distribution noise.
Example 1 Example 2
$L_{2}$ distance RPE FPR FNR $L_{2}$ distance RPE FPR FNR
(I) 20/20/400 Ridge 8.191(0.722) 7.175(2.487) 0.967(0.012) 0.000(0.000) 8.836(0.616) 7.774(2.098) 0.966(0.014) 0.001(0.009)
Lasso 12.751(1.752) 4.356(3.484) 0.094(0.023) 0.504(0.113) 11.259(1.153) 6.126(3.125) 0.093(0.023) 0.501(0.124)
Alasso 12.334(1.762) 3.310(2.229) 0.082(0.020) 0.436(0.105) 11.172(1.484) 5.296(2.938) 0.082(0.022) 0.445(0.130)
Enet 6.907(1.215) 5.497(2.813) 0.430(0.043) 0.007(0.020) 8.110(1.035) 6.943(2.481) 0.430(0.050) 0.033(0.066)
SRL 12.658(1.636) 3.998(2.627) 0.104(0.026) 0.467(0.110) 11.234(1.495) 5.686(3.090) 0.104(0.026) 0.444(0.130)
SRIG-o 8.765(0.672) 13.523(2.362) 0.569(0.060) 0.000(0.000) 9.310(0.524) 12.251(1.736) 0.566(0.053) 0.009(0.040)
SRIG 8.810(0.662) 13.683(2.361) 0.571(0.060) 0.000(0.000) 9.232(0.539) 12.053(1.737) 0.564(0.054) 0.009(0.040)
GSRE-o 3.528(1.223) 1.827(2.352) 0.023(0.022) 0.000(0.000) 4.714(1.402) 2.668(2.794) 0.026(0.022) 0.013(0.066)
GSRE 3.416(1.272) 1.861(2.440) 0.024(0.022) 0.000(0.000) 4.621(1.372) 2.558(2.753) 0.026(0.023) 0.013(0.066)
(II) 40/40/400 Ridge 6.985(0.612) 3.544(0.911) 0.962(0.014) 0.000(0.000) 7.345(0.594) 3.850(1.053) 0.964(0.013) 0.000(0.000)
Lasso 9.290(1.360) 1.565(0.671) 0.057(0.020) 0.263(0.086) 7.612(1.374) 2.354(1.268) 0.044(0.017) 0.195(0.101)
Alasso 9.445(1.394) 1.035(0.441) 0.050(0.021) 0.235(0.090) 7.703(1.568) 1.582(0.840) 0.037(0.014) 0.172(0.088)
Enet 5.373(0.719) 4.123(1.389) 0.308(0.052) 0.000(0.000) 5.968(0.773) 4.158(1.312) 0.321(0.047) 0.005(0.018)
SRL 9.194(1.349) 1.459(0.635) 0.067(0.028) 0.240(0.089) 7.446(1.459) 2.229(1.221) 0.052(0.024) 0.172(0.099)
SRIG-o 3.323(1.039) 1.761(1.494) 0.032(0.051) 0.000(0.000) 4.244(1.466) 2.514(2.179) 0.064(0.085) 0.000(0.000)
SRIG 3.149(1.073) 1.730(1.455) 0.031(0.053) 0.000(0.000) 3.957(1.307) 2.135(1.852) 0.050(0.073) 0.000(0.000)
GSRE-o 2.455(0.514) 0.640(0.448) 0.014(0.012) 0.000(0.000) 3.014(0.616) 0.806(0.529) 0.015(0.014) 0.000(0.000)
GSRE 2.370(0.544) 0.652(0.476) 0.013(0.010) 0.000(0.000) 3.193(0.785) 1.109(0.908) 0.014(0.012) 0.000(0.000)
(III) 60/60/400 Ridge 6.196(0.733) 2.211(0.599) 0.961(0.015) 0.000(0.000) 6.350(0.748) 2.267(0.624) 0.961(0.014) 0.000(0.000)
Lasso 7.443(1.239) 0.930(0.403) 0.035(0.016) 0.121(0.073) 5.743(1.231) 1.253(0.730) 0.023(0.014) 0.048(0.065)
Alasso 8.456(1.316) 0.658(0.223) 0.035(0.016) 0.167(0.069) 6.482(1.164) 0.961(0.423) 0.023(0.014) 0.080(0.067)
Enet 4.679(0.742) 3.452(1.313) 0.210(0.060) 0.000(0.000) 5.316(0.985) 3.687(1.664) 0.203(0.057) 0.001(0.009)
SRL 7.900(1.302) 0.991(0.561) 0.262(0.095) 0.103(0.085) 6.216(1.283) 1.347(0.951) 0.198(0.133) 0.029(0.080)
SRIG-o 2.482(0.407) 0.759(0.360) 0.004(0.007) 0.000(0.000) 2.774(0.615) 0.813(0.540) 0.002(0.006) 0.000(0.000)
SRIG 2.266(0.428) 0.724(0.370) 0.004(0.007) 0.000(0.000) 2.664(0.628) 0.764(0.528) 0.003(0.006) 0.000(0.000)
GSRE-o 2.109(0.320) 0.358(0.193) 0.014(0.014) 0.000(0.000) 2.434(0.491) 0.450(0.287) 0.007(0.009) 0.000(0.000)
GSRE 2.017(0.324) 0.350(0.187) 0.013(0.014) 0.000(0.000) 2.408(0.480) 0.438(0.263) 0.007(0.010) 0.000(0.000)
@@
Table 3: Performance comparison of estimation, prediction and model selection for Example 1&2 with laplace distribution noise.
Example 1 Example 2
$L_{2}$ distance RPE FPR FNR $L_{2}$ distance RPE FPR FNR
(I) 20/20/400 Ridge 8.091(0.739) 7.074(2.383) 0.966(0.014) 0.000(0.000) 8.830(0.575) 7.685(1.974) 0.968(0.015) 0.001(0.009)
Lasso 12.317(1.589) 4.254(3.995) 0.088(0.022) 0.484(0.120) 11.483(1.619) 6.117(3.243) 0.092(0.019) 0.476(0.113)
Alasso 12.032(1.558) 3.380(3.020) 0.078(0.022) 0.425(0.117) 11.089(1.610) 5.076(2.782) 0.081(0.023) 0.419(0.112)
Enet 6.765(1.326) 5.553(3.003) 0.431(0.055) 0.008(0.029) 8.064(0.996) 6.587(2.399) 0.437(0.044) 0.040(0.079)
SRL 12.239(1.581) 3.724(3.096) 0.110(0.028) 0.464(0.113) 11.358(1.755) 5.339(2.379) 0.104(0.029) 0.437(0.105)
SRIG-o 8.818(0.639) 13.685(2.280) 0.567(0.069) 0.007(0.047) 9.276(0.545) 12.158(1.759) 0.568(0.062) 0.005(0.023)
SRIG 8.869(0.640) 13.866(2.309) 0.568(0.069) 0.007(0.047) 9.205(0.540) 11.978(1.728) 0.567(0.062) 0.005(0.023)
GSRE-o 3.233(0.932) 1.363(1.051) 0.026(0.032) 0.000(0.000) 4.437(1.515) 2.242(2.907) 0.026(0.019) 0.015(0.066)
GSRE 3.120(0.903) 1.375(1.067) 0.027(0.035) 0.000(0.000) 4.320(1.453) 2.098(2.780) 0.026(0.018) 0.013(0.066)
(II) 40/40/400 Ridge 6.801(0.584) 3.635(1.211) 0.963(0.018) 0.000(0.000) 7.351(0.585) 3.835(0.905) 0.965(0.013) 0.000(0.000)
Lasso 8.939(1.370) 1.641(0.960) 0.050(0.021) 0.227(0.103) 8.017(1.273) 2.531(1.459) 0.045(0.019) 0.203(0.092)
Alasso 8.954(1.544) 0.996(0.624) 0.039(0.017) 0.193(0.087) 8.252(1.275) 1.863(1.105) 0.039(0.017) 0.187(0.087)
Enet 5.379(0.826) 4.206(1.560) 0.300(0.046) 0.000(0.000) 5.993(0.821) 4.110(1.471) 0.331(0.053) 0.000(0.000)
SRL 8.661(1.369) 1.534(0.889) 0.057(0.027) 0.205(0.097) 7.801(1.271) 2.379(1.411) 0.054(0.025) 0.191(0.097)
SRIG-o 3.239(0.974) 1.641(1.506) 0.023(0.052) 0.000(0.000) 4.340(1.520) 2.643(2.235) 0.075(0.090) 0.000(0.000)
SRIG 3.039(1.135) 1.644(1.723) 0.022(0.054) 0.000(0.000) 4.208(1.469) 2.496(2.108) 0.072(0.087) 0.000(0.000)
GSRE-o 2.455(0.502) 0.664(0.498) 0.011(0.014) 0.000(0.000) 3.077(0.744) 0.812(0.634) 0.015(0.017) 0.000(0.000)
GSRE 2.360(0.550) 0.678(0.515) 0.010(0.013) 0.000(0.000) 3.040(0.738) 0.803(0.622) 0.016(0.017) 0.000(0.000)
(III) 60/60/400 Ridge 6.041(0.730) 2.095(0.518) 0.963(0.012) 0.000(0.000) 6.419(0.693) 2.326(0.660) 0.961(0.013) 0.000(0.000)
Lasso 7.127(1.236) 0.920(0.357) 0.031(0.015) 0.115(0.080) 6.000(0.996) 1.332(0.648) 0.028(0.016) 0.067(0.065)
Alasso 7.888(1.435) 0.606(0.279) 0.032(0.018) 0.143(0.090) 6.589(1.155) 0.972(0.391) 0.024(0.016) 0.087(0.068)
Enet 4.727(0.803) 3.612(1.374) 0.197(0.054) 0.000(0.000) 5.402(0.896) 3.800(1.379) 0.196(0.046) 0.000(0.000)
SRL 7.069(1.254) 0.879(0.357) 0.036(0.018) 0.103(0.084) 5.812(1.041) 1.223(0.573) 0.033(0.021) 0.053(0.056)
SRIG-o 2.489(0.423) 0.781(0.344) 0.003(0.006) 0.000(0.000) 2.875(0.534) 0.856(0.463) 0.004(0.007) 0.000(0.000)
SRIG 2.281(0.401) 0.748(0.322) 0.003(0.006) 0.000(0.000) 2.772(0.549) 0.808(0.448) 0.004(0.007) 0.000(0.000)
GSRE-o 2.127(0.356) 0.400(0.201) 0.010(0.012) 0.000(0.000) 2.574(0.419) 0.505(0.269) 0.008(0.010) 0.000(0.000)
GSRE 2.047(0.337) 0.394(0.196) 0.010(0.012) 0.000(0.000) 2.544(0.420) 0.504(0.265) 0.008(0.010) 0.000(0.000)
@@
Table 4: Performance comparison of estimation, prediction and model selection for Example 1&2 with uniform distribution noise.
Example 1 Example 2
$L_{2}$ distance RPE FPR FNR $L_{2}$ distance RPE FPR FNR
(I) 20/20/400 Ridge 8.130(0.745) 7.266(2.614) 0.968(0.015) 0.001(0.009) 8.830(0.588) 7.580(2.125) 0.971(0.013) 0.007(0.024)
Lasso 12.460(1.422) 4.432(4.073) 0.092(0.023) 0.485(0.121) 11.528(1.547) 6.079(3.241) 0.092(0.021) 0.481(0.130)
Alasso 12.401(1.626) 3.546(3.249) 0.086(0.023) 0.467(0.126) 11.187(1.704) 5.231(3.177) 0.079(0.024) 0.433(0.131)
Enet 6.852(1.315) 5.624(3.094) 0.431(0.054) 0.011(0.037) 8.073(1.028) 6.625(2.472) 0.433(0.047) 0.043(0.076)
SRL 12.744(1.382) 3.982(3.170) 0.110(0.026) 0.472(0.113) 11.317(1.631) 5.299(2.578) 0.104(0.030) 0.444(0.119)
SRIG-o 8.816(0.652) 13.680(2.313) 0.570(0.072) 0.007(0.047) 9.287(0.545) 12.187(1.757) 0.563(0.067) 0.005(0.023)
SRIG 8.859(0.658) 13.837(2.352) 0.572(0.069) 0.007(0.047) 9.219(0.545) 12.016(1.741) 0.560(0.067) 0.005(0.023)
GSRE-o 3.491(1.099) 1.755(1.540) 0.021(0.026) 0.000(0.000) 4.451(1.435) 2.222(2.600) 0.030(0.023) 0.013(0.066)
GSRE 3.337(1.051) 1.674(1.471) 0.025(0.030) 0.000(0.000) 4.397(1.424) 2.194(2.586) 0.028(0.024) 0.013(0.066)
(II) 40/40/400 Ridge 6.979(0.531) 3.544(1.024) 0.969(0.010) 0.000(0.000) 7.356(0.488) 3.661(0.884) 0.967(0.012) 0.000(0.000)
Lasso 9.248(1.365) 1.712(0.923) 0.057(0.020) 0.252(0.097) 7.915(1.254) 2.397(1.164) 0.045(0.019) 0.201(0.095)
Alasso 9.360(1.335) 1.061(0.540) 0.043(0.017) 0.209(0.077) 8.142(1.318) 1.750(1.076) 0.040(0.016) 0.189(0.075)
Enet 5.418(0.767) 4.217(1.431) 0.306(0.046) 0.000(0.000) 5.965(0.763) 4.127(1.346) 0.327(0.049) 0.000(0.000)
SRL 9.095(1.487) 1.566(0.863) 0.063(0.026) 0.233(0.090) 7.658(1.193) 2.223(1.228) 0.048(0.020) 0.176(0.086)
SRIG-o 3.415(1.105) 1.903(1.751) 0.030(0.053) 0.000(0.000) 4.347(1.428) 2.617(2.068) 0.076(0.096) 0.000(0.000)
SRIG 3.170(1.175) 1.804(1.767) 0.029(0.053) 0.000(0.000) 4.117(1.320) 2.325(1.877) 0.064(0.089) 0.000(0.000)
GSRE-o 2.522(0.537) 0.712(0.553) 0.012(0.013) 0.000(0.000) 2.998(0.616) 0.738(0.530) 0.013(0.015) 0.000(0.000)
GSRE 2.426(0.572) 0.704(0.564) 0.013(0.014) 0.000(0.000) 2.952(0.598) 0.722(0.512) 0.012(0.015) 0.000(0.000)
(III) 60/60/400 Ridge 6.190(0.657) 2.164(0.515) 0.960(0.014) 0.000(0.000) 6.359(0.695) 2.335(0.731) 0.963(0.010) 0.000(0.000)
Lasso 7.456(1.232) 0.972(0.341) 0.036(0.017) 0.132(0.080) 5.922(1.064) 1.294(0.707) 0.028(0.016) 0.068(0.065)
Alasso 8.237(1.195) 0.644(0.224) 0.036(0.015) 0.168(0.074) 6.580(1.087) 0.990(0.482) 0.023(0.012) 0.092(0.070)
Enet 4.754(0.783) 3.651(1.340) 0.196(0.052) 0.000(0.000) 5.354(1.007) 3.756(1.610) 0.199(0.053) 0.001(0.009)
SRL 7.362(1.143) 0.913(0.306) 0.040(0.020) 0.121(0.078) 5.809(1.026) 1.251(0.623) 0.030(0.018) 0.059(0.057)
SRIG-o 2.503(0.429) 0.782(0.336) 0.003(0.005) 0.000(0.000) 2.874(0.565) 0.863(0.521) 0.004(0.007) 0.000(0.000)
SRIG 2.311(0.399) 0.755(0.321) 0.003(0.005) 0.000(0.000) 2.761(0.575) 0.805(0.498) 0.004(0.008) 0.000(0.000)
GSRE-o 2.152(0.366) 0.410(0.181) 0.010(0.010) 0.000(0.000) 2.540(0.416) 0.472(0.271) 0.009(0.012) 0.000(0.000)
GSRE 2.082(0.348) 0.402(0.184) 0.011(0.012) 0.000(0.000) 2.507(0.426) 0.464(0.270) 0.009(0.011) 0.000(0.000)
Tables 5-8 display the performance of each method for Example 3 under four different types of noise. Similarly, as the sample size increases, the numerical performance of each algorithm improves. Ridge regression shows a low FNR but a high FPR, which indicate a tendency towards overfitting. Lasso and Alasso methods generally have lower values of FPR, but they have higher values for $L_{2}$ distance and FNR. In all four challenging environments, the GSRE method consistently outperforms other methods in model estimation and prediction, which shows its robust performance. By balancing the values of FPR and FNR, it is clear that our method also performs very well in model selection for Example 3.
Table 5: Performance comparison of estimation, prediction and model selection for Example 3 with normal distribution noise.
$L_{2}$ distance RPE FPR FNR
(I) 20/20/400 Ridge 9.229 (0.688) 7.278 (1.696) 0.963 (0.016) 0.005 (0.018)
Lasso 11.060 (1.308) 6.902 (2.401) 0.093 (0.020) 0.508 (0.125)
Alasso 10.931 (1.282) 6.352 (2.383) 0.084 (0.023) 0.468 (0.127)
Enet 8.831 (0.971) 6.781 (1.958) 0.412 (0.043) 0.111 (0.105)
SRL 10.872 (1.471) 6.471 (2.510) 0.171 (0.027) 0.392 (0.150)
SRIG-o 10.396 (0.383) 11.385 (1.172) 0.846 (0.073) 0.020 (0.051)
SRIG 9.811 (0.514) 9.960 (1.353) 0.912 (0.041) 0.011 (0.031)
GSRE-o 7.039 (1.580) 4.199 (2.615) 0.288 (0.109) 0.052 (0.110)
GSRE 6.145 (1.487) 3.121 (2.296) 0.321 (0.118) 0.025 (0.080)
(II) 40/40/400 Ridge 7.667 (0.563) 3.791 (1.027) 0.960 (0.017) 0.000 (0.000)
Lasso 7.668 (1.052) 2.859 (1.093) 0.049 (0.018) 0.197 (0.087)
Alasso 7.825 (1.161) 2.195 (0.919) 0.043 (0.016) 0.205 (0.079)
Enet 6.959 (0.752) 4.416 (1.170) 0.306 (0.048) 0.019 (0.040)
SRL 7.513 (1.127) 2.674 (1.052) 0.056 (0.025) 0.176 (0.087)
SRIG-o 10.263 (0.281) 11.230 (1.061) 0.495 (0.095) 0.009 (0.027)
SRIG 9.583 (0.440) 9.706 (1.208) 0.589 (0.117) 0.001 (0.009)
GSRE-o 5.631 (1.318) 2.874 (1.971) 0.040 (0.037) 0.013 (0.040)
GSRE 4.861 (1.354) 2.074 (1.915) 0.031 (0.022) 0.007 (0.034)
(III) 60/60/400 Ridge 6.839 (0.687) 2.325 (0.511) 0.959 (0.011) 0.000 (0.000)
Lasso 5.825 (1.062) 1.628 (0.720) 0.044 (0.022) 0.073 (0.075)
Alasso 6.159 (1.073) 1.208 (0.458) 0.026 (0.017) 0.091 (0.076)
Enet 6.821 (0.895) 4.644 (1.253) 0.190 (0.056) 0.012 (0.032)
SRL 5.738 (1.101) 1.625 (0.750) 0.042 (0.025) 0.068 (0.072)
SRIG-o 4.973 (0.896) 2.253 (0.958) 0.027 (0.017) 0.000 (0.000)
SRIG 4.207 (0.789) 1.468 (0.705) 0.042 (0.019) 0.000 (0.000)
GSRE-o 4.093 (0.957) 1.301 (0.784) 0.029 (0.024) 0.001 (0.009)
GSRE 3.835 (0.814) 1.124 (0.633) 0.026 (0.018) 0.000 (0.000)

### v2

@@ 4.1 Simulation data
Table 1: Performance comparison of estimation, prediction and model selection for Example 3 with Gaussian noise.
$L_{2}$ distance RPE FPR FNR MCC
(I) 20/20/400 Ridge 8.181(1.355) 5.224(2.651) 0.948(0.022) 0.010(0.027) 0.071(0.056)
Lasso 10.790(1.225) 6.233(2.605) 0.072(0.019) 0.489(0.142) 0.452(0.142)
Alasso 10.963(1.113) 6.352(2.476) 0.051(0.018) 0.509(0.148) 0.484(0.153)
Enet 8.000(1.268) 3.221(1.878) 0.138(0.027) 0.152(0.108) 0.591(0.105)
SRL 11.259(1.259) 6.894(2.434) 0.116(0.029) 0.481(0.124) 0.379(0.126)
SRIG-o 10.393(0.359) 11.403(1.097) 0.852(0.060) 0.019(0.049) 0.139(0.051)
SRIG 10.393(0.359) 11.403(1.097) 0.852(0.060) 0.019(0.049) 0.139(0.051)
GSRE-o 7.095(1.540) 4.613(2.541) 0.318(0.105) 0.031(0.074) 0.483(0.125)
GSRE 7.095(1.540) 4.613(2.541) 0.318(0.105) 0.031(0.074) 0.483(0.125)
(II) 40/40/400 Ridge 7.755(0.560) 3.800(1.064) 0.954(0.017) 0.000(0.000) 0.083(0.016)
Lasso 8.186(1.035) 3.724(1.461) 0.028(0.018) 0.280(0.107) 0.732(0.096)
Alasso 8.039(1.238) 2.137(0.898) 0.034(0.024) 0.225(0.085) 0.756(0.091)
Enet 6.827(0.945) 3.940(1.385) 0.118(0.037) 0.059(0.071) 0.692(0.079)
SRL 8.236(1.174) 3.898(1.517) 0.028(0.018) 0.289(0.121) 0.726(0.108)
SRIG-o 7.649(1.432) 6.121(2.508) 0.335(0.200) 0.001(0.009) 0.531(0.220)
SRIG 7.649(1.432) 6.121(2.508) 0.335(0.200) 0.001(0.009) 0.531(0.220)
GSRE-o 5.145(1.057) 2.129(1.316) 0.039(0.033) 0.013(0.038) 0.887(0.079)
GSRE 5.145(1.057) 2.129(1.316) 0.039(0.033) 0.013(0.038) 0.887(0.079)
(III) 60/60/400 Ridge 6.819(0.620) 2.354(0.521) 0.957(0.011) 0.000(0.000) 0.081(0.011)
Lasso 6.615(1.597) 2.682(1.693) 0.019(0.017) 0.180(0.160) 0.829(0.096)
Alasso 6.796(1.099) 1.310(0.557) 0.026(0.019) 0.141(0.082) 0.833(0.082)
Enet 6.074(1.148) 3.389(1.448) 0.068(0.031) 0.037(0.062) 0.801(0.065)
SRL 6.061(1.247) 2.011(1.082) 0.029(0.022) 0.105(0.110) 0.849(0.087)
SRIG-o 4.911(0.862) 2.182(0.918) 0.025(0.018) 0.000(0.000) 0.926(0.045)
SRIG 4.911(0.862) 2.182(0.918) 0.025(0.018) 0.000(0.000) 0.926(0.045)
GSRE-o 4.266(0.914) 1.454(0.731) 0.023(0.023) 0.001(0.009) 0.934(0.057)
GSRE 4.266(0.914) 1.454(0.731) 0.023(0.023) 0.001(0.009) 0.934(0.057)
@@
Table 2: Performance comparison of estimation, prediction and model selection for Example 3 with T-distribution noise.
$L_{2}$ distance RPE FPR FNR MCC
(I) 20/20/400 Ridge 7.503(1.441) 4.224(2.498) 0.953(0.017) 0.001(0.009) 0.082(0.024)
Lasso 9.299(2.075) 4.521(2.763) 0.062(0.022) 0.375(0.196) 0.564(0.177)
Alasso 10.301(1.703) 5.345(2.686) 0.044(0.020) 0.457(0.176) 0.547(0.161)
Enet 6.492(2.081) 2.348(2.108) 0.116(0.027) 0.113(0.129) 0.653(0.129)
SRL 10.548(1.437) 5.809(2.439) 0.113(0.021) 0.413(0.133) 0.435(0.129)
SRIG-o 10.370(0.355) 11.352(1.084) 0.842(0.062) 0.023(0.050) 0.141(0.049)
SRIG 10.370(0.355) 11.352(1.084) 0.842(0.062) 0.023(0.050) 0.141(0.049)
GSRE-o 6.067(2.006) 3.474(2.631) 0.308(0.096) 0.027(0.063) 0.494(0.122)
GSRE 6.067(2.006) 3.474(2.631) 0.308(0.096) 0.027(0.063) 0.494(0.122)
(II) 40/40/400 Ridge 7.048(0.941) 3.185(1.268) 0.951(0.013) 0.000(0.000) 0.087(0.012)
Lasso 5.718(2.606) 2.039(2.176) 0.040(0.035) 0.140(0.166) 0.798(0.130)
Alasso 4.648(3.015) 1.051(2.080) 0.026(0.037) 0.081(0.129) 0.878(0.146)
Enet 5.550(1.828) 2.768(2.040) 0.119(0.053) 0.043(0.073) 0.705(0.098)
SRL 5.911(2.682) 2.231(2.324) 0.040(0.035) 0.148(0.177) 0.794(0.144)
SRIG-o 6.748(1.884) 4.865(2.962) 0.235(0.216) 0.000(0.000) 0.650(0.246)
SRIG 6.748(1.884) 4.865(2.962) 0.235(0.216) 0.000(0.000) 0.650(0.246)
GSRE-o 3.463(1.554) 1.067(1.625) 0.046(0.034) 0.007(0.034) 0.872(0.085)
GSRE 3.463(1.554) 1.067(1.625) 0.046(0.034) 0.007(0.034) 0.872(0.085)
(III) 60/60/400 Ridge 5.677(0.966) 1.624(0.804) 0.958(0.012) 0.000(0.000) 0.080(0.012)
Lasso 3.497(1.866) 0.744(1.205) 0.043(0.029) 0.025(0.093) 0.866(0.106)
Alasso 3.378(4.223) 0.834(3.931) 0.018(0.049) 0.027(0.087) 0.938(0.135)
Enet 3.995(1.654) 1.479(1.530) 0.108(0.044) 0.015(0.049) 0.742(0.086)
SRL 3.528(2.249) 0.790(1.540) 0.047(0.048) 0.028(0.093) 0.861(0.134)
SRIG-o 4.421(1.037) 1.801(1.175) 0.032(0.059) 0.000(0.000) 0.919(0.081)
SRIG 4.421(1.037) 1.801(1.175) 0.032(0.059) 0.000(0.000) 0.919(0.081)
GSRE-o 2.720(1.205) 0.618(0.960) 0.033(0.035) 0.001(0.009) 0.911(0.081)
GSRE 2.720(1.205) 0.618(0.960) 0.033(0.035) 0.001(0.009) 0.911(0.081)
@@
Table 3: Performance comparison of estimation, prediction and model selection for Example 3 with Laplace noise.
$L_{2}$ distance RPE FPR FNR MCC
(I) 20/20/400 Ridge 8.072(1.420) 5.017(2.618) 0.953(0.018) 0.011(0.025) 0.065(0.050)
Lasso 10.352(1.464) 5.791(2.887) 0.071(0.021) 0.456(0.177) 0.480(0.178)
Alasso 11.112(1.386) 6.363(2.460) 0.050(0.022) 0.531(0.141) 0.471(0.168)
Enet 7.837(1.450) 3.158(1.804) 0.133(0.027) 0.155(0.113) 0.596(0.113)
SRL 10.967(1.373) 6.487(2.826) 0.118(0.033) 0.475(0.134) 0.382(0.145)
SRIG-o 10.390(0.337) 11.386(1.029) 0.839(0.060) 0.021(0.049) 0.144(0.055)
SRIG 10.390(0.337) 11.386(1.029) 0.839(0.060) 0.021(0.049) 0.144(0.055)
GSRE-o 6.915(1.644) 4.334(2.494) 0.309(0.105) 0.027(0.069) 0.493(0.102)
GSRE 6.915(1.644) 4.334(2.494) 0.309(0.105) 0.027(0.069) 0.493(0.102)
(II) 40/40/400 Ridge 7.653(0.645) 3.759(1.161) 0.952(0.016) 0.000(0.000) 0.085(0.015)
Lasso 7.975(1.209) 3.643(1.610) 0.028(0.018) 0.265(0.124) 0.743(0.104)
Alasso 7.553(1.284) 1.896(0.861) 0.029(0.020) 0.204(0.111) 0.782(0.113)
Enet 6.653(0.994) 3.791(1.404) 0.115(0.029) 0.060(0.066) 0.693(0.078)
SRL 8.001(1.266) 3.875(1.831) 0.026(0.020) 0.284(0.143) 0.737(0.110)
SRIG-o 7.675(1.516) 6.179(2.520) 0.338(0.197) 0.003(0.019) 0.522(0.211)
SRIG 7.675(1.516) 6.179(2.520) 0.338(0.197) 0.003(0.019) 0.522(0.211)
GSRE-o 4.919(1.207) 2.006(1.384) 0.038(0.035) 0.005(0.023) 0.895(0.081)
GSRE 4.919(1.207) 2.006(1.384) 0.038(0.035) 0.005(0.023) 0.895(0.081)
(III) 60/60/400 Ridge 6.677(0.751) 2.267(0.614) 0.955(0.015) 0.000(0.000) 0.083(0.015)
Lasso 6.310(1.666) 2.446(1.717) 0.020(0.016) 0.160(0.149) 0.838(0.078)
Alasso 6.047(1.358) 1.053(0.505) 0.024(0.019) 0.105(0.076) 0.862(0.080)
Enet 6.131(1.201) 3.518(1.546) 0.057(0.026) 0.043(0.060) 0.822(0.051)
SRL 5.790(1.234) 1.804(0.917) 0.028(0.017) 0.085(0.101) 0.864(0.075)
SRIG-o 4.854(0.701) 2.101(0.734) 0.020(0.013) 0.000(0.000) 0.939(0.036)
SRIG 4.854(0.701) 2.101(0.734) 0.020(0.013) 0.000(0.000) 0.939(0.036)
GSRE-o 3.947(0.799) 1.213(0.627) 0.021(0.019) 0.001(0.009) 0.939(0.050)
GSRE 3.947(0.799) 1.213(0.627) 0.021(0.019) 0.001(0.009) 0.939(0.050)